THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 93 — 1 June 2008

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Editorial

Dear Colleagues,

It is my pleasure to present you the 93^{rd} issue of the Magellanic Clouds Newsletter, featuring the rich variety of topics that are uniquely studied in these intriguing galaxies. We must indeed cherish our luck to have these celestial bodies so near us (and not behind the Galactic Centre!).

Today we also say goodbye to co-editor Snežana Stanimirović, whose sanity, wisdom, and above all great friendship, have always been crucial support — including in reviving this newsletter. Thanks heaps Snežana, we’ll surely see more of you amongst The Clouds!

We look forward to seeing you all at IAU Symposium 256, ”The Magellanic System: Stars, Gas, and Galaxies”, at Keele University from 28 July until 1 August. The registration remains open until 18 July, and poster abstracts can be submitted up to that date (further oral contributions may be considered in exceptional circumstances). Posters will receive ample attention including a special poster viewing session and brief oral introductions to selected posters, and all posters enter a competition (with prizes!). There is also the opportunity to have your spectacular images displayed in large format (of order ten square metres) — if you are interested in this possibility then please contact the organisers at iaus256@astro.keele.ac.uk. More details can be found in the announcements that are posted on the website http://www.astro.keele.ac.uk/iaus256.

The next issue will be distributed on the 2^{nd} of August 2008; the deadline for contributions is the 1^{st} of August.

Editorially Yours, Jacco van Loon

Refereed Journal Papers

An optical & X-ray study of the counterpart to the SMC X-ray binary pulsar system SXP 327

M. J. Coe ^{1}, M. Schurch^{1}, V.A. McBride^{1}, L.J. Townsend^{1}, R. Corbet^{2}, J. Galache^{3} and A. Udalski^{4}

1University of Southampton, UK 2GSFC/NASA, USA 3CfA, USA 4Warsaw, Poland

Optical and X-ray observations are presented here of a newly reported X-ray transient system in the Small Magellanic Cloud. The data reveal many previously unknown X-ray detections of this system and clear evidence for a 49.995d binary period. In addition, the optical photometry show recurring outburst features at the binary period which may well be indicative of the neutron star interacting with a circumstellar disk around a Be star.

Accepted for publication in MNRAS Available from arXiv:0803.3941

The Initial Mass Function of the Stellar Association NGC 602 in the Small Magellanic Cloud with Hubble Space Telescope ACS Observations

Markus Schmalzl^{1}, Dimitrios A. Gouliermis^{1}, Andrew E. Dolphin^{2} and Thomas Henning^{1}

1Max Planck Institute for Astronomy, Heidelberg, Germany 2Raytheon Corporation, MA, USA

We present our photometric study of the stellar association NGC 602 in the wing of the Small Magellanic Cloud (SMC). The data were taken in the filters F555W and F814W using the Advanced Camera for Surveys (ACS) on-board the Hubble Space Telescope (HST). Photometry was performed using the ACS module of the stellar photometry package DOLPHOT. We detected more than 5,500 stars with a magnitude range of 14< ∼m_{555}< ∼28 mag. Three prominent stellar concentrations are identified with star counts in the observed field, the association NGC 602 itself, and two clusters, one of them not being currently in any known catalog. The Color-Magnitude Diagrams (CMDs) of both clusters show features typical for young open clusters, while that of the association reveals bright main sequence (MS) and faint pre-main sequence (PMS) stars as the members of the system. We construct the initial mass spectrum (IMS) of the association by applying an age-independent method of counting the PMS stars within evolutionary tracks, while for the bright MS stars we transform their magnitudes to masses with the use of mass-luminosity relations. The IMS of NGC 602 is found to be well represented by a single-power law, corresponding to an Initial Mass Function (IMF) of slope Γ ≈ −1.2 for 1< ∼M/M_{⊙}< ∼45. This indicates that the shape of the IMF of a star forming system in the SMC for stars with masses higher than 1 M_{⊙} seems to be quite similar to the field IMF in the solar neighborhood.

Accepted for publication in ApJ Available from arXiv:0804.0543 and from http://www.astro.uni-bonn.de/∼dgoulier/Science/NGC602/ms.pdf

The chemical composition of red giant stars in four intermediate-age clusters of the Large Magellanic Cloud

A. Mucciarelli^{1}, E. Carretta^{2}, L. Origlia^{2} and F. R. Ferraro^{1}

1Dipartimento di Astronomia, Bologna, Italy 2INAF-Osservatorio Astronomico di Bologna, Italy

This paper presents the chemical abundance analysis of a sample of 27 red giant stars located in 4 populous intermediateage globular clusters in the Large Magellanic Cloud, namely NGC 1651, 1783, 1978 and 2173. This analysis is based on high-resolution (R ∼ 47, 000) spectra obtained with the UVES@VLT spectrograph. For each cluster we derived up to 20 abundance ratios sampling the main chemical elemental groups, namely light odd-Z, alpha, iron-peak and neutroncapture elements. All the analysed abundance patterns behave similarly in the 4 clusters and also show negligible star-to-star scatter within each cluster. We find [Fe/H]= −0.30 ± 0.03, −0.35 ± 0.02, −0.38 ± 0.02 and −0.51 ± 0.03 dex for NGC 1651, 1783, 1978 and 2173, respectively. The measurement of light odd-Z nuclei gives slightly subsolar [Na/Fe] and a more significant [Al/Fe] depletion (∼ −0.50 dex). The [α/Fe] abundance ratios are nearly solar, while the iron-peak elements well trace that one of the iron. s-process elements behave in a peculiar way: light s-elements give subsolar [Y/Fe] and [Zr/Fe] abundance ratios, while heavy s-elements give enhanced [Ba/Fe], [La/Fe] and [Nd/Fe] with respect to the solar values. Also, the [Eu/Fe] abundance ratio turns out to be enhanced (∼ 0.4 dex).

Accepted for publication in Astronomical Journal Available from arXiv:0804.4061

Spatially resolved dusty torus toward the red supergiant WOH G64 in the Large Magellanic Cloud

Keiichi Ohnaka^{1}, Thomas Driebe^{1}, Karl-Heinz Hofmann^{1}, Gerd Weigelt^{1} and Markus Wittkowski^{2}

1Max-Planck-Institut für Radioastronomie, Germany 2European Southern Observatory, Germany

We present N-band spectro-interferometric observations of the red supergiant WOH G64 in the Large Magellanic Cloud (LMC) using MIDI at the Very Large Telescope Interferometer (VLTI). While the very high luminosity (∼ 5 × 10^{5} L_{⊙}) previously estimated for WOH G64 suggests that it is a very massive star with an initial mass of ∼40 M_{⊙}, its low effective temperature (∼3200 K) is in serious disagreement with the current stellar evolution theory. WOH G64 was observed with VLTI/MIDI using the UT2-UT3 and UT3-UT4 baseline configurations. The dust envelope around WOH G64 has been spatially resolved with a baseline of ∼60 m — the first MIDI observations to resolve an individual stellar source in an extragalactic system. The observed N-band visibilities show a slight decrease from 8 to ∼10 μm and a gradual increase longward of ∼10 μm, reflecting the 10 μm silicate feature in selfabsorption. This translates into a steep increase of the uniform-disk diameter from 8 to 10 μm (from 18 to 26 mas) and a roughly constant diameter above 10 μm. The visibilities measured at four position angles differing by ∼60^{∘} but at approximately the same baseline length (∼60 m) do not show a noticeable difference, suggesting that the object appears nearly centrosymmetric. The observed N-band visibilities and spectral energy distribution can be reproduced by an optically and geometrically thick silicate torus model viewed close to pole-on. The luminosity of the central star is derived to be ∼2.8 × 10^{5} L_{⊙}, which is by a factor of 2 lower than the previous estimates based on spherical models. We also identify the H_{2}O absorption features at 2.7 and 6 μm in the spectra obtained with the Infrared Space Observatory and the Spitzer Space Telescope. The 2.7 μm feature originates in the photosphere and/or the extended molecular layers, while the 6 μm feature is likely to be of circumstellar origin. The lower luminosity newly derived from our MIDI observations and two-dimensional modeling brings the location of WOH G64 on the H-R diagram in much better agreement with theoretical evolutionary tracks for a 25 M_{⊙} star. However, the effective temperature is still somewhat too cool for the theory. The low effective temperature of WOH G64 places it very close to or even beyond the Hayashi limit, which implies that this object may be experiencing unstable, violent mass loss.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:0803.3823

Characteristic Scales in Stellar Clustering: A Transition Near the Disk Scale Height

Mary Crone Odekon^{1}

1Skidmore College, USA

The autocorrelation function provides an objective test for the existence of special scales in the hierarchical clustering of young stars. We apply this measure to single-star photometry for the brightest main sequence stars in the Small Magellanic Cloud (SMC), the Large Magellanic Cloud (LMC), M 33, and M 31, using data from the Magellanic Clouds Photometric Survey and the Massey Local Group Survey. Our primary result is the identification of a transition to a higher correlation dimension (weaker clustering) at one kpc in the LMC and M 31, and at 300 pc in M 33. We suggest that this transition marks the large-scale regime where disk geometry and dynamics set the scale for structure. On smaller scales, the correlation functions for each galaxy are scale-free over at least two orders of magnitude, with a projected correlation dimension varying from 1.0 for M 31 to 1.8 for the SMC. This variation is probably caused by a combination of differences in stellar ages and masses, physical environment, and extinction.

Accepted for publication in The Astrophysical Journal Available from arXiv:0804.4607

New XMM-Newton observations of SNRs in the SMC

M.D. Filipović^{1,2}, F. Haberl^{1}, P.F. Winkler^{3}, W. Pietsch^{1}, J.L. Payne^{4}, E.J. Crawford^{2}, A.Y. De Horta^{2}, F.H. Stootman^{2} and B.E. Reaser^{5}

1Max-Planck-Institut für extraterrestrische Physik, Germany 2University of Western Sydney, Australia 3Department of Physics, Middlebury College, USA 4James Cook University, Australia 5Swarthmore College, USA

A complete overview of the supernova remnant (SNR) population is required to investigate their evolution and interaction with the surrounding interstellar medium in the Small Magellanic Cloud (SMC). Recent XMM-Newton observations of the SMC cover three known SNRs (DEM S5, SNR B0050−72.8, and SNR B0058−71.8), which are poorly studied and are X-ray faint. We used new multi-frequency radio-continuum surveys and new optical observations at Hα, S ii, and O iii wavelengths, in combination with the X-ray data, to investigate their properties and to search for new SNRs in the SMC. We used X-ray source selection criteria and found one SMC object with typical SNR characteristics (HFPK 334), that was initially detected by ROSAT. We analysed the X-ray spectra and present multi-wavelength morphological studies of the three SNRs and the new candidate. Using a non-equilibrium ionisation collisional plasma model, we find temperatures kT around 0.18 keV for the three known remnants and 0.69 keV for the candidate. The low temperature, low surface brightness, and large extent of the three remnants indicates relatively large ages. The emission from the new candidate (HFPK 334) is more centrally peaked and the higher temperature suggests a younger remnant. Our new radio images indicate that a pulsar wind nebulae (PWN) is possibly associated with this object. The SNRs known in the SMC show a variety of morphological structures that are relatively uncorrelated in the different wavelength bands, probably caused by the different conditions in the surrounding medium with which the remnant interacts.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:0805.0165

Optical Spectra of Radio Planetary Nebulae in the Small Magellanic Cloud

J. L. Payne^{1}, M. D. Filipović^{2}, E. J. Crawford^{2}, A.Y. De Horta^{2}, G. L. White^{1} and F.H. Stootman^{2}

1Centre for Astronomy, James Cook University, Australia 2University of Western Sydney, Australia

We present preliminary results from spectral observations of four candidate radio sources co-identified with known planetary nebulae (PNe) in the Small Magellanic Cloud (SMC). These were made using the Radcliffe 1.9-meter telescope in Sutherland, South Africa. These radio PNe were originally found in Australia Telescope Compact Array (ATCA) surveys of the SMC at 1.42 and 2.37 GHz, and were further confirmed by new high resolution ATCA images at 6 and 3 cm (4′′/2′′). Optical PNe and radio candidates are within 2′′ and may represent a subpopulation of selected radio bright objects. Nebular ionized masses of these objects may be 2.6 M_{⊙} or greater, supporting the existence of PNe progenitor central stars with masses up to 8 M_{⊙}.

Accepted for publication in SAJ Available from arXiv:0804.4539

Radio Continuum Study of Supernova Remnants in the Large Magellanic Cloud — SNR J0519−6926

E.J. Crawford^{1}, M.D. Filipović^{1} and J.L. Payne^{2}

1University of Western Sydney, Australia 2James Cook University, Australia

We present the results of new high resolution ATCA observations of SNR J0519−6926. We found that this SNR exhibits a typical ”horseshoe” appearance with α = −0.55 ± 0.08 and D = 28 ± 1 pc. No polarization (or magnetic fields) are detected to a level of 1%. This is probably due to a relatively poor sampling of the uv plane caused be observing in ”snap-shot” mode.

Accepted for publication in SAJ Available from arXiv:0804.4537

Structures in surface-brightness profiles of LMC and SMC star clusters: evidence of mergers?

Luziane Carvalho^{1}, Tiago A. Saurin^{1}, Eduardo Bica^{1}, Charles J. Bonatto^{1} and Alex A. Schmidt^{2}

1IF/UFRGS, Brasil 2UFSM, Brasil

The LMC and SMC are rich in binary star clusters, and some mergers are expected. It is thus important to characterise single clusters, binary clusters and candidates for mergers. We selected a sample of star clusters in each Cloud with this aim. Surface photometry of 25 SMC and 22 LMC star clusters was carried with the ESO Danish 1.54 m telescope. 23 clusters were observed for the first time for these purposes. We fitted Elson, Fall and Freeman (EFF) profiles to the data, deriving structural parameters, luminosities and masses. We also use isophotal maps to constrain candidates for cluster interactions. The structural parameters, luminosities and masses presented good agreement with those in the literature. Three binary clusters in the sample have a double profile. Four clusters (NGC 376, K 50, K 54 and NGC 1810) do not have companions and present important deviations from EFF profiles. The present sample contains blue and red Magellanic clusters. Profiles with excess with respect to EFF were detected in some blue clusters. We find evidence that important deviations from the body of EFF profiles might be used as a tool to detect cluster mergers.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:0804.4110

The Araucaria Project. The Distance of the Large Magellanic Cloud from Near-Infrared Photometry of RR Lyrae Variables

Olaf Szewczyk^{1,2}, Grzegorz Pietrzyński^{1,2}, Wolfgang Gieren^{1}, Jesper Storm^{3}, Alistair Walker^{4}, Luca Rizzi^{5}, Karen Kinemuchi^{1,6}, Fabio Bresolin^{7}, Rolf-Peter Kudritzki^{7} and Massimo Dall’Ora^{8}

1Universidad de Concepción, Departamento de Fisica, Astronomy Group, Casilla 160-C, Concepción, Chile 2Warsaw University Observatory, Al. Ujazdowskie 4, 00-478, Warsaw, Poland 3Astrophysikalisches Institut Potsdam, An der Sternwarte 16, D-14482 Potsdam, Germany 4Cerro Tololo Inter-American Observatory, Casilla 603, La Serena, Chile 5Joint Astronomy Centre, 66 N. A’ohoku Pl., Hilo, Hawaii, 96720, USA 6University of Florida, Department of Astronomy, Gainesville, Florida 32611-2055, USA 7Institute for Astronomy, University of Hawaii at Manoa, 2680 Woodlawn Drive, Honolulu HI 96822, USA 8INAF, Osservatorio Astronomico di Capodimonte, I-80131 Napoli, Italy

We have obtained deep infrared J and K band observations of five fields located in the Large Magellanic Cloud (LMC) bar with the ESO New Technology Telescope equipped with the SOFI infrared camera. In our fields, 65 RR Lyrae stars catalogued by the OGLE collaboration were identified. Using different theoretical and empirical calibrations of the period-luminosity-metallicity relation, we find consistent LMC distance moduli values. Since the observed fields are situated very close to the center of the LMC, the correction for the tilt of the LMC bar with respect to the line of sight is negligible. Our adopted best true distance modulus to the LMC of 18.58±0.03 (statistical) ± 0.11 (systematic) mag agrees very well with most independent determinations to this galaxy.

Accepted for publication in AJ Available from arXiv:0804.3333

Formation of the SMC: an acient major merger as a solution for the kinematical difference between old stars and H i gas

Kenji Bekki^{1}

1UNSW, Australia

Recent observations of the Small Magellanic Cloud (SMC) have revealed that the H i gas shows a significant amount of rotation (V_{c} ∼ 60 km s^{−1}), while no or little rotation is evident for the old stellar populations. We suggest that this unique kinematical difference between these components in the SMC can be caused by a major merger event which occurred in the early stage of the SMC formation. Our simulations show that dissipative dwarf-dwarf merging can transform two gas-rich dwarf irregulars into a new dwarf, which consists of a spheroidal stellar component and a rotating extended H i disk. The remnant of this dwarf-dwarf merging shows significantly different kinematics between stars and gas, in the sense that a gas disk rotates rapidly while a stellar component shows little rotation. We thus suggest that the simulated dwarf having a dynamically hot spheroid and an extended gas disk finally evolves into the present SMC after efficient stripping of the outer gas via tidal fields of the Galaxy and the Large Magellanic Cloud. We also suggest that spatial distributions and kinematics of RGB and AGB stars with different ages in the possible spheroidal component of the SMC can provide valuable information on whether and when a past major merger event really occurred in the SMC.

Accepted for publication in ApJL Available from arXiv:0804.4563

The scattered debris of the Magellanic Stream

Tobias Westmeier^{1} and Bärbel S. Koribalski^{1}

1Australia Telescope National Facility, Australia

Searching the H i Parkes All-Sky Survey (HIPASS) and its northern extension, we detected a population of very compact high-velocity clouds (HVCs) with similar velocities in the Galactic standard-of-rest frame which appear to be arranged in several filaments aligned with the nearby Magellanic Stream. A comparison with published O vi/Ca ii absorption and H i emission line measurements suggests that the HVCs are condensations within an extended and mainly ionised component of the Magellanic Stream. They coincide in position with a faint gas stream predicted in numerical simulations of the Magellanic Clouds by Gardiner & Noguchi (1996). Consequently, the Magellanic Stream could be much more extended than generally believed.

Accepted for publication in MNRAS Available from arXiv:0805.0820

The Molecular Ridge Close to 30 Doradus in the Large Magellanic Cloud

Jürgen Ott^{1,2}, Tony Wong^{3}, Jorge L. Pineda^{4}, Annie Hughes^{5}, Erik Muller^{6}, Zhi-Yun Li^{7}, Min Wang^{8}, Lister Staveley-Smith^{9}, Yasuo Fukui^{10}, Axel Weiß^{11}, Christian Henkel^{11} and Ulrich Klein^{4}

1National Radio Astronomy Observatory, USA 2California Institute of Technology, USA 3University of Illinois, USA 4University of Bonn, Germany 5Swinburne University of Technology, Australia 6CSIRO Australia Telescope National Facility, Australia 7University of Virginia, USA 8Purple Mountain Observatories, China 9University of Western Australia, Australia 10Nagoya University, Japan 11Max-Planck-Institut für Radioastronomie, Germany

With the ATNF Mopra telescope we are performing a survey in the ^{12}CO(1-0) line to map the molecular gas in the Large Magellanic Cloud (LMC). For some regions we also obtained interferometric maps of the high density gas tracers HCO^{+} and HCN with the Australia Telescope Compact Array (ATCA). Here we discuss the properties of the elongated molecular complex that stretches about 2 kpc southward from 30 Doradus. Our data suggests that the complex, which we refer to as the “molecular ridge,” is not a coherent feature but consists of many smaller clumps that share the same formation history. Likely molecular cloud formation triggers are shocks and shearing forces that are present in the surrounding south-eastern H i overdensity region, a region influenced by strong ram pressure and tidal forces. The molecular ridge is at the western edge of the the overdensity region where a bifurcated velocity structure transitions into a single disk velocity component. We find that the ^{12}CO(1-0) and H i emission peaks in the molecular ridge are typically near each other but never coincide. A likely explanation is the conversion of warmer, low-opacity H i to colder, high-opacity H i from which H_{2} subsequently forms. On smaller scales, we find that very dense molecular gas, as traced by interferometric HCO^{+} and HCN maps, is associated with star formation along shocked filaments and with rims of expanding shell-like structures, both created by feedback from massive stars.

Accepted for publication in Publications of the Astronomical Society of Australia Available from arXiv:0805.1452

Spitzer Mid-infrared Study of Compact H ii Regions in the Magellanic Clouds

V. Charmandaris^{1,2}, M. Heydari-Malayeri^{2} and E. Chatzopoulos^{1}

1University of Crete, Greece 2Observatoire de Paris, France

We present a study of the mid-infrared properties and dust content of a sample of 27 H ii “blobs”, a rare class of compact H ii regions in the Magellanic Clouds. A unique feature of this sample is that even though these H ii regions are of high and low excitation they have nearly the same physical sizes ∼ 1.5 − 3 pc. We base our analysis on archival 3 − 8 μm infrared imagery obtained with the Infrared Array Camera (IRAC) on board the Spitzer Space Telescope. We find that despite their youth, sub-solar metallicity and varied degrees of excitation, the mid-infrared colors of these regions are similar to those of typical H ii regions. Higher excitation “blobs” (HEBs) display stronger 8 μm emission and redder colors than their low-excitation counterparts (LEBs).

Accepted for publication in A&A Available from arXiv:0805.2337 and from http://users.physics.uoc.gr/∼vassilis/papers/blobs spitzer.pdf

The Hα Galaxy Survey VI. Star-forming companions of nearby field galaxies

P. A. James^{1}, J. O’Neill^{2} and N. S. Shane^{3}

1Liverpool JMU, UK 2Wirral Grammar School for Girls, UK 3MSSL, UK

We present a search for star-forming satellite galaxies that are close enough to their parent galaxies to be considered analogues of the Magellanic Clouds. Our search technique relied on the detection of the satellites in continuumsubtracted narrow-band Hα imaging of the central galaxies, which removes most of the background and foreground line-of-sight companions, thus giving a high probability that we are detecting true satellites. The search was performed for 119 central galaxies at distances between 20 and 40 Mpc, although spatial incompleteness means that we have effectively searched 53 full satellite-containing volumes. We find only 9 probable star-forming satellites, around 9 different central galaxies, and 2 possible satellites. After incompleteness correction, this is equivalent to 0.17/0.21 satellites per central galaxy. The Small Magellanic Cloud is just below the median values of both star formation rate and R-band luminosity of the 9 probable satellites. The Large Magellanic Cloud, however, has a higher R-band luminosity than any of the 9 and is only exceeded in star formation rate by the one satellite that appears to be undergoing a tidally-induced starburst. Thus the Milky Way appears to be quite unusual, both in having two starforming satellite galaxies and in the high luminosity of the Large Magellanic Cloud.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:0805.2817

A New Outburst In the Extraordinary Central Star of LMC-N66

Miriam Peña^{1}, Maria Teresa Ruiz^{2}, Patricio Rojo^{2}, Silvia Torres-Peimbert^{1} and Wolf-Rainer Hamann^{3}

1Instituto de Astronomía, UNAM, México 2Departamento de Astronomía, Universidad de Chile, Chile 3Universität Potsdam, Germany

This is the first report on the new outburst presented by the central star of the LMC-N66 nebula. This object was classified as a planetary nebula, however, its true nature is under debate. In the period 1955–1990 the central star was almost undetectable and only nebular emission lines were observed. In 1990, the beginning of an outburst was detected and in few months it became much brighter and developed wide He and N lines, typical of a Wolf Rayet star of the N-sequence. The maximum occurred in 1994 and afterwards the star slowly faded. Analysis of its evolution showed that it has a variable mass-loss rate which occasionally increases enormously, creating a false photosphere at a much larger radius, making it appear a few magnitudes brighter. The present outburst has occurred 13 years after the episode from 1994 to 2000. So far this new event has similar characteristics although there are some significant differences in the spectral features. We present optical and FUSE spectra showing the main properties of this latter event.

Accepted for publication in ApJ Letters Available from arXiv:0805.2184

Cepheids in external galaxies. I. The maser-host galaxy NGC 4258 and the metallicity dependence of P-L and P-W relations

Bono, G.^{1,2}, Caputo, F.^{1}, Fiorentino, G.^{3}, Marconi, M.^{4} and Musella, I.^{4}

1INAF-Osservatorio Astronomico di Roma, Via Frascati 33, 00040, Monte Porzio Catone, Italy 2European Southern Observatory, Karl-Schwarzschild-Str. 2, 85748 Garching bei München, Germany 3Kapteyn Astronomical Institute, University of Groningen, Postbus 800, 9700 AV Groningen, The Netherlands 4INAF-Osservatorio Astronomico di Capodimonte, Via Moiariello 16, 80131 Napoli, Italy

We perform a detailed analysis of Cepheids in NGC 4258, Magellanic Clouds and Milky Way in order to verify the reliability of the theoretical scenario based on a large set of nonlinear convective pulsation models. We derive Wesenheit functions from the synthetic BVI magnitudes of the pulsators and we show that the sign and the extent of the metallicity effect on the predicted Period-Wesenheit (P-W) relations change according to the adopted passbands. These P-W relations are applied to measured BVI magnitudes of NGC 4258, Magellanic and Galactic Cepheids available in the literature. We find that Magellanic and Galactic Cepheids agree with the metallicity dependence of the predicted P-W relations. Concerning the NGC 4258 Cepheids, the results strongly depend on the adopted metallicity gradient across the galactic disc. The most recent nebular oxygen abundances support a shallower gradient and provide a metallicity dependence that agrees well with current pulsation predictions. Moreover, the comparison of Cepheid distances based on VI magnitudes with distance estimates based on the revised TRGB method for external galaxies, on the HST trigonometric parallaxes for Galactic Cepheids, and on eclipsing binaries in the Magellanic Clouds seems to favor the metallicity correction predicted by pulsation models. The sign and the extent of the metallicity dependence of the Period-Wesenheit and of the Period-Luminosity relations change according to the adopted passbands. Therefore, distances based on different methods and/or bands should not be averaged. The use of extragalactic Cepheids to constrain the metallicity effect requires new accurate and extensive nebular oxygen measurements.

Accepted for publication in ApJ Available from arXiv:0805.1592

AGB stars of the intermediate-age LMC cluster NGC 1846. II. Dredge up along the AGB

T. Lebzelter^{1}, M. T. Lederer^{1}, S. Cristallo^{2}, K.H. Hinkle^{3}, O. Straniero^{2} and B. Aringer^{1,4}

1Department of Astronomy, University of Vienna, Austria 2INAF, Osservatorio Astronomico di Collurania, Italy 3NOAO, Tucson, USA 4Dipartimento di Astronomia, Universita di Padova, Italy

Aims: We investigate the change in the surface abundance of ^{12}C during the evolution along the AGB, aiming to constrain third dredge-up models.

Methods: High-resolution, near-infrared spectra of a sample of AGB stars in the LMC cluster NGC 1846 were obtained. A cluster sample ensures a high level of homogeneity with respect to age, metallicity, and distance. The C/O ratio and the ratio of ^{12}C/^{13}C were measured and compared with our evolutionary models. Results: For the first time, we show the evolution of the C/O and ^{12}C/^{13}C ratios along a cluster AGB. Our findings allow us to check the reliability of the evolutionary models and, in particular, the efficiency of the third dredge up. The increase in both C/O and ^{12}C/^{13}C in the observed O-rich stars is reproduced by the models well. However, the low carbon isotopic ratios of the two C-stars in our sample indicate the late occurrence of moderate extra mixing. The extra mixing affects the most luminous AGB stars and is capable of increasing the abundance of ^{13}C, while leaving unchanged the C/O ratio, which has been fixed by the cumulative action of several third dredge-up episodes. We find indications that the F abundance also increases along the AGB, supporting an in situ production of this element.

Accepted for publication in A&A Available from arXiv:0805.3242